OpenStudy (anonymous):

If f(x) = x + 4, find f(3).

OpenStudy (baldymcgee6):

If f(x) = x + 4, find f(3) f(3) = 3 +4

OpenStudy (baldymcgee6):

Simply sub the value of x into the function

OpenStudy (anonymous):

im soo confused what does it mean by find f(3) if it already shows f(x)?

OpenStudy (unklerhaukus):

\(f(x)\) means that the value of the dependent variable \(f\) is a function of the independent variable \(x\) , the value of \(f\) depends of the value of \(x\) \(f(3)\) is the function evaluated at \(x=3\)

OpenStudy (unklerhaukus):